# OpenVoice: Versatile Instant Voice Cloning

### 03 Dec 2023 v1

@@ 2.3 Training
Second, we require flow layers to eliminate as much tone color information as possible from the audio features. During training, for each audio sample, its text is converted to a sequence of phonemes in International Phonetic Alphabet [1], and each phoneme is represented by a learnable vector embedding. The sequence of vector embedding is passed to a transformer [15] encoder to produce the feature representation of the text content. Denote this feature as $\mathbf{L}\in\mathbb{R}^{c\times l}$, where $c$ is the number of feature channels and $l$ is the number of phonemes in the input text. The audio waveform is processed by the encoder and flow layers to produce the feature representation $\mathbf{Z}\in\mathbb{R}^{c\times t}$, where $t$ is the length of the features along the time dimension. Then we align $\mathbf{L}$ with $\mathbf{Z}$ along the time dimension using dynamic time warping [13, 10] (an alternative is monotonic alignment [5, 6]) to produce $\mathbf{\bar{L}}\in\mathbb{R}^{c\times t}$, and minimize the KL-divergence between $\mathbf{\bar{L}}$ and $\mathbf{Z}$. Since $\mathbf{\bar{L}}$ does not contain any tone color information, the minimization objective would encourage the flow layers to remove tone color information from their output $\mathbf{Z}$. The flow layers are conditioned on the tone color information from the tone color encoder, which further helps the flow layers to identify what information needs to be eliminated. In addition, we do not provide any style or language information for the flow layers to condition on, which prevents the flow layers to eliminate information other than tone color. Since the flow layers are invertible, conditioning them on a new piece of tone color information and running its inverse process can add the new tone color back to the feature representations, which are then decoded to the raw waveform with the new tone color embodied.

### 13 Dec 2023 v2

@@ 2.3 Training
Second, we require flow layers to eliminate as much tone color information as possible from the audio features. During training, for each audio sample, its text is converted to a sequence of phonemes in International Phonetic Alphabet (IPA) [1], and each phoneme is represented by a learnable vector embedding. The sequence of vector embedding is passed to a transformer [15] encoder to produce the feature representation of the text content. Denote this feature as $\mathbf{L}\in\mathbb{R}^{c\times l}$, where $c$ is the number of feature channels and $l$ is the number of phonemes in the input text. The audio waveform is processed by the encoder and flow layers to produce the feature representation $\mathbf{Z}\in\mathbb{R}^{c\times t}$, where $t$ is the length of the features along the time dimension. Then we align $\mathbf{L}$ with $\mathbf{Z}$ along the time dimension using dynamic time warping [13, 10] (an alternative is monotonic alignment [5, 6]) to produce $\mathbf{\bar{L}}\in\mathbb{R}^{c\times t}$, and minimize the KL-divergence between $\mathbf{\bar{L}}$ and $\mathbf{Z}$. Since $\mathbf{\bar{L}}$ does not contain any tone color information, the minimization objective would encourage the flow layers to remove tone color information from their output $\mathbf{Z}$. The flow layers are conditioned on the tone color information from the tone color encoder, which further helps the flow layers to identify what information needs to be eliminated. In addition, we do not provide any style or language information for the flow layers to condition on, which prevents the flow layers to eliminate information other than tone color. Since the flow layers are invertible, conditioning them on a new piece of tone color information and running its inverse process can add the new tone color back to the feature representations, which are then decoded to the raw waveform with the new tone color embodied.
@@ 3 Experiment
Importance of IPA. We found that using IPA as the phoneme dictionary is crucial for the tone color converter to perform cross-lingual voice cloning. As we detailed in Section 2.3, in training the tone color converter, the text is first converted into a sequence of phonemes in IPA, then each phoneme is represented by a learnable vector embedding. The sequence of embedding is encoded with transformer layers and compute loss against the output of the flow layers, aiming to eliminate the tone color information. IPA itself is a cross-lingual unified phoneme dictionary, which enables the flow layers to produce a language-neutral representation. Even if we input a speech audio with unseen language to the tone color converter, it is still able to smoothly process the audio. We also experimented with other types of phoneme dictionaries but the resulting tone color converter tend to mispronounce some phonemes in unseen languages. Although the input audio is correct, there is a high likelihood that the output audio is problematic and sounds non-native.